OpenStudy (anonymous):

I am a little confused about what my textbook is saying about this problem. Could someone help please? The temperature outside is 10° Celsius. You would like to know how many degrees this is in Fahrenheit. You know the formula for degrees Celsius is: C = 5 over 9 (F − 32). Part 1: Solve the equation for F and show all steps. Part 2: Determine what 10 degrees Celsius is equivalent to in Fahrenheit.

OpenStudy (anonymous):

\[C=\frac{5}{9}\left(F-32\right)\] You have C as a function of F and want you to get F as a function of C so solve for F(C). Then calculate with that equation F(10) is.

OpenStudy (anonymous):

Ok that make a little bit more since... so what you are saying is that i need to use that problem to find the answer to part 1? how would I calculate the equation F(10)? would it be like 10 = 5/9 (F - 32) ?

OpenStudy (anonymous):

You could do it that way and solve for F. But they are walking you through manipulating the first equation, which is C(F) = (5/9)(F-32) and want you to change it so F is alone on one side with numbers and C on the other. Then you just have to plug in 10 for C and do the math. For example, C(32) = (5/9)(32-32) = (5/9)(0)=0.

OpenStudy (shamim):

\[we know the equation, c \div5=(f-32)/9 or, 9*c/5=f-32 or, f=(9*c/5+32)or, f=9*10/5+32 or, f=50\]

hero (hero):

\[C=\frac{5}{9}\left(F-32\right) \\ C=\frac{5\left(F-32\right)}{9} \\ C=\frac{5F-160}{9} \\ 9C=5F-160\\ 9C + 160=5F\\ \frac{9C + 160}{5}=F\\ F = \frac{9}{5}C + 32\]

hero (hero):

Now, let C = 10 to get: \[F = \frac{9}{5}(10) + 32\] Now, solve for F

hero (hero):

Should get something like F = 18 + 32 F = 50

OpenStudy (anonymous):

Ok,i think i understand, but is this answering question 1 or 2? sorry that i'm so slow at this....

OpenStudy (shamim):

i think hero is right

OpenStudy (anonymous):

Question 1 is the solving the equation for F and leaving C as is. Question 2 is replacing C with 10 and then calculating the value for F. Don't forget to use units in your answer.

OpenStudy (anonymous):

Thanks everyone!